OpenStudy (anonymous):

7(x - 3) + 3(4 - x) = -10

OpenStudy (anonymous):

4x-9=-10

OpenStudy (anonymous):

If your question is solve for x then just multiply 7 with the terms in parenthesis and 3 with the terms in parenthesis and add like terms together. Then solve for x

OpenStudy (anonymous):

Distribute and then solve for x. 7x - 21 + 12 - 3x = -10 4x - 9 = -10 4x = -1 x = -1/4

OpenStudy (anonymous):

\[7x-21+12-3x=-10\] Now solve for x